OpenStudy (anonymous):

arclength of s(t)=(cos t, sin t, cos 2t, sin 2t) for t in [0, pi]

OpenStudy (zarkon):

compute \[\int\limits_{0}^{\pi}\|s'(t)\|dt\]

OpenStudy (zarkon):

I get \(\pi\sqrt{5}\)

OpenStudy (anonymous):

Awesome, it seemed so simple that I thought that I did something wrong